A harmonic oscillator vibrates with frequency of 4 HZ and an amplitude of 2cm.what is its maximum velocity?

1
Expert's answer
2021-12-17T11:50:04-0500

"V_{max} = Aω \\

ω = 2 \pi f = 8 \pi \;rad/s \\

A = 2 \;cm = 0.02 \; m\\

V_{max} = 0.02 \times 8 \pi = 0.5 \;m/s = 16 \; \pi \; cm/s"